OpenStudy (anonymous):

A sporting goods store sells right-handed and left-handed baseball gloves. In one month, 15 gloves were sold for a total revenue of $702. Right-handed gloves cost $42 and left-handed gloves cost $51. How many left-handed gloves did they sell?

OpenStudy (anonymous):

if they sell say \(x\) right handed gloves, then since the total is \(15\) they must have solve \(15-x\) left handed gloves then you can solve \[42x+51(15-x)=702\]for \(x\)

OpenStudy (ybarrap):

You are being asked for the number of left-handed gloves, @satellite73 let x be the number of right-handed gloves. So once you solve for x, use 15-x to find the number of left-handed gloves. Was that what tripped you here @michelle.k ?

OpenStudy (anonymous):

and when I want to solve for the right hand I use this formula 42(15-x)+51=702?

OpenStudy (anonymous):

missing an \(x\) but close \[42(15-x)+51x=702\]

OpenStudy (ybarrap):

There! @satellite73 made it a little more clear. One equation is necessary, that's all. Now x is the number of left-handed gloves.

OpenStudy (anonymous):

@satellite73 I got 7 but they said that's wrong

OpenStudy (ybarrap):

$$ 42(15-x)+51x=702\\ x(51-42)=702-42\times15\\ x=\cfrac{702-42\times 15}{51-42}=8 $$